OpenStudy (anonymous):

How do you factor and simplify the algebraic expression:

OpenStudy (anonymous):

\[(x+9)^{1/4}+(x+9)^{3/4}\]

OpenStudy (valpey):

\[\text{Let }A = (x+9)^{1/4}\]Then you just have to factor:\[A+A^3\]

OpenStudy (anonymous):

What happened to the fractions?

OpenStudy (valpey):

If \[A = Z^{1/4}\]Then \[A^3 = Z^{3/4}\]

OpenStudy (anonymous):

I don't understand

OpenStudy (valpey):

Okay, do you agree that:\[x^{1/4}*x^{1/4}*x^{1/4}=x^{3/4}\]

OpenStudy (anonymous):

YES

OpenStudy (anonymous):

Oops, caps, sorry about that.

OpenStudy (precal):

do you have your final answer?

OpenStudy (anonymous):

No, I don't understand how to find it.

OpenStudy (precal):